if the hypotenuse of a right triangle is 15cm long and one leg is 9 how long is the other leg?
Answer by richwmiller(17219)

a^2+b^2=c^2 where and b are the legs of the right triangle and c is the hypotenuse
substitute 15 for c and 9 for a and solve for b
9^2+b^2=15^2